A medical doctor threw a small medical package horizontally at 10m/s from the 5th floor of the hospital 15m high. How long does the medical package take to reach the ground floor?


1
Expert's answer
2021-02-04T19:09:36-0500
"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot15\ m}{9.8\dfrac{m}{s^2}}}=1.75\ s."